The probability distribution shown represents the number of trips of five nights or more that American adults take per year. (That is, 6% do not take any trips lasting five nights or more, 70% take one trip lasting five nights or more per year, etc.) Find the mean.


1
Expert's answer
2021-12-16T17:17:39-0500
"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:}
 Number & 0 & 1 & 2 & 3 & 4 \\ 
 of\ trips \\ 
\hline
 P(X=x) & 0.06 & 0.70 & 0.20 & 0.03& 0.01 \\
 
\end{array}"

"mean=E(X)=\sum_ix_iP(X=x_i)"

"=0(0.06)+1(0.70)+2(0.20)+3(0.03)+4(0.01)"

"=1.23"

The mean is "1.23."